OpenStudy (anonymous):

Please Help ! FACTOR THE POLYNOMIAL: A sports team is building a new stadium on a rectangular lot of land. If the lot measures 6x by 10x and the sports field will be 1x by 4x, how much of the lot will be left over to build bleachers on?

OpenStudy (anonymous):

The area of the lot is \[6x \times 10x=60x^2\]The area of the sports field is\[x \times 4x = 4x^2\]The area left over is (area of lot) - (area of sports field) =\[60x^2-4x^2=56x^2\]

OpenStudy (anonymous):

Duude! You Rock, Thanks Soo Much. . . I Was just About To Post a similar Question But Now I Understand Sorta, Lol, Thanks!

OpenStudy (anonymous):

np ;]